Recent Log Entries

COMMENT ON Replace MQA1C17 Trim Card

Log Number: 3829989Logbook(s): Body: 

In Situ trim card unresponsive to set point after reseating. Replaced module, set current and cycled successfully. Handed back to Ops. All is well again.

dogCalc V0-10 download

Log Number: 3829978Logbook(s): Tags: Body: dogCalc V0-10 download

CED workspace: OPS
CED login_key: pro

CHANNELNEW VALUEOLD VALUEMDOG9E.BDL-1085564.7187-1105230.4513MBZ9E02.BDL1080134.71871099800.4513

Checking 1st pass separation

Log Number: 3829973Logbook(s): Entry Makers: mockBody: 

Needed a slight adjustment of Vernier, phase, and 2S10H corrector to get BPM positions nominal.
Vernier: 0.74 -> 0.71
Phase: 39 -> 40
2S10H corrector -> +7 Bdl

Follow-up Re: Re-check Hall B noise after Injector steering

Log Number: 3829970Logbook(s): Body: 

Nice sleuthing. Both Wiens are between 1I02 and 0I02, so four power supplies for the four electrodes, maybe explains similar but separate ~10 kHz ferquencies. Note, we're replacing all those supplies this SAD. An option for further troubleshooting is to set WIens for 0 deg, turn the HVPS OFF, and re-measure beam motion.